A student purchases the following lock. There are 40 numbers. The three number key cannot use the same number twice. How many possible 3 digit keys are possible for this lock?


1
Expert's answer
2021-02-02T02:24:32-0500

"\boxed{A_k^n=\dfrac {k!} {(k-n)!}}"


We are choosing n = 3 out of k = 40


A = "\dfrac {37!*38*39*40}{37!} = 59280"


Answer: 59280 possible keys